OpenStudy (anonymous):

if tan= 9/12, what is cos

OpenStudy (anonymous):

tan=9/12 (we know that these are the exact dimensions, b/c otherwise the would be minimized.) So, Adj=12 Opp=9 Hyp=\[\sqrt{9^{2}+12^{2}}\]\[\sqrt{81+144}\]\[\sqrt{225}\]\[±15\] the distance, or the side, can't be = to a negative, so leave it as hyp=15 . So, Cos= adj / hyp = 12/15 (or you can, in fact even better, say 4/5 )